# Revisiting Multimodal Positional Encoding in Vision–Language Models

### v1

@@ 1 Introduction
The permutation-invariant nature of the self-attention mechanism requires the use of positional encodings to inform Large Language Models (LLMs) of sequence order, relative distance, and structural dependencies. While early methods relied on absolute position embeddings (Vaswani et al., 2017), relative encodings—which better generalize to varying sequence lengths—have become the standard. Among these, Rotary Position Embedding (RoPE) (Su et al., 2024) has emerged as the de facto choice in modern LLMs such as Llama (Grattafiori et al., 2024) and Qwen (Yang et al., 2025).
Vision-Language Models (VLMs) also require positional encodings that can handle heterogeneous modalities, including 1D text and 2D/3D visual inputs. Current methods fall into two main categories: 1D sequential and multi-dimensional designs. The former, exemplified by vanilla RoPE (Su et al., 2024) and V2PE (Ge et al., 2024), flattens and concatenates all inputs into a single sequence. While simple, this approach discards the native visual geometry, leading to a significant degradation in performance on tasks requiring visual grounding and spatial reasoning.
Multi-dimensional designs, the second approach, extend RoPE to multiple axes (time, height, width) by partitioning embedding channels. Qwen2‑VL Wang et al. (2024a) adopts Multimodal RoPE (MRoPE) to unify positional encoding for text and visual tokens. However, MRoPE allocates the position embedding into t-h-w chunk, placing the temporal infomation entirely in the high‑frequency channels. This bias in temporal encoding harms long‑range video modeling. Subsequent work has attempted to improve it, but this has led to a fragmented landscape of highly specialized solutions. Some methods focus exclusively on image understanding (Wang et al., 2025), others on video comprehension (Wei et al., 2025; Li et al., 2025; Liu et al., 2025), and a third group on image generation (Liao et al., 2025; Wu et al., 2025). While these models achieve notable performance in their respective domains, the development of a truly robust and versatile VLM requires a more holistic positional encoding strategy. In this work, we aim to develop a more holistic positional encoding strategy capable of supporting the core, unified capabilities of image and video understanding, complemented by fine-grained visual grounding.
@@ 2.1 Preliminaries: Vanilla RoPE
Vanilla RoPE (Su et al., 2024) is a pivotal method for encoding positional information in modern LLMs. Unlike additive position embeddings, RoPE applies a rotational transformation to the query and key vectors, thereby incorporating relative position dependencies directly into the self-attention mechanism. Given a query vector $\bm{q}$ at position $m$ and a key vector $\bm{k}$ at position $n$, the attention scores $\bm{S}$ are calculated as:
@@ 2.2 Position Design
The most straightforward approach, employed by vanilla RoPE (Su et al., 2024) and V2PE (Ge et al., 2024), is to treat the multimodal input as a flattened, one-dimensional sequence. Position indices are assigned incrementally, with the position $m_{i}$ of the $i$-th token defined as $m_{i}=m_{i-1}+s_{\text{mod}}$, where $s_{\text{mod}}$ is a step size specific to the token’s modality. For vanilla RoPE, all modalities are treated uniformly, with $s=1$.
As shown in Figure 1(a), this design presents two significant drawbacks. First, it discards the inherent 3D structure of visual content, which can alter the spatio-temporal reasoning capabilities of a VLM. Second, position indices can grow exceedingly large in long sequences, negatively affecting the model’s extrapolation performance (Wei et al., 2025).
To address the issue of large position indices, V2PE (Ge et al., 2024) introduces dynamic position scaling for visual tokens, setting their step size $s_{\text{visual}}$ to a value in $\{1,1/2,\dots,1/256\}$. This modification mitigates the rapid growth of position indices and has shown benefits in long video understanding. However, the 3D structure of visual content is still ignored in 1D sequential design.
@@
To preserve the native 3D structure of visual content, methods like MRoPE (Wang et al., 2024a) (See Figure 1(b)) extend the scalar position identifier to a multi-dimensional tuple. For instance, a token’s position can be represented as $\bm{m}_{i}=(m_{i}^{t},m_{i}^{h},m_{i}^{w})$, corresponding to its temporal, vertical, and horizontal axes. MRoPE conceptually treats each visual content (e.g., an image or a set of video frames) as a single, large ”cube”. The temporal position of the subsequent token is then set by ”jumping” past the maximum coordinate value of current block. This is achieved with the update rule:
@@
This strategy guarantees that no positional overlaps occur between modalities. However, proponents of VideoRoPE (Li et al., 2025) and HoPE (Li et al., 2025) argue that MRoPE’s position design lacks inter-modal symmetry, they introduce a ”diagonal layout” by centering the spatial coordinates (see Figure 1(c)). In this scheme, visual frames are not only stacked along the temporal axis but are also shifted along the vertical and horizontal axes. Despite its theoretical elegance, this diagonal layout introduces a critical flaw: the potential for position id overlap between visual content and generated text tokens. For high-resolution image content like documents, the spatial coordinates of visual tokens can extend into the index range subsequently assigned to the generated text tokens. We identify this positional ambiguity as a source of ”modalities confusion in generation”, a failure mode that manifested as endless text repetition in our later experiments.
CircleRoPE (Wang et al., 2025) arranges image tokens in a circular layout, orthogonal to the linear axis of text positions (see Figure 1(d)). A key property of this design is that it renders all visual tokens equidistant from any given text token, which theoretically promotes uniform attention across the image. However, CircleRoPE’s design has two limitations. First, the large interval between modalities may impede effective cross-modal interaction. Second, lacking a temporal axis, it collapses all video frames onto a single ring, which introduces severe temporal ambiguity.
@@ 2.3.3 Towards an Optimal Frequency Allocation
Multi-Head Allocation. The first strategy, termed Multi-Head RoPE (MHRoPE), is inspired by recent work demonstrating channel-level redundancy in RoPE (e.g., partial RoPE (Barbero et al., 2025)). Based on the premise that similar redundancy exists at the attention head level, MHRoPE partitions the positional encoding task among different attention heads [^4] For Group Query Attention (GQA), we partition on KV heads and repeat on corresponding query heads., as shown in Figure 3. A primary advantage of this strategy is that each axis is encoded using the full frequency spectrum available within its assigned heads. This approach avoids the loss of frequency resolution inherent to channel-splitting methods. Moreover, it may be more scalable. As the number of positional axes grows (Liu et al., 2025), partitioning a fixed channel channels (e.g., 128) becomes untenable, whereas dedicating distinct heads to new dimensions offers a far more robust and flexible approach.
Interleaved Allocation. Our second strategy, employed in our MRoPE-Interleaved (MRoPE-I) method, distributes feature channels to the temporal ($t$), vertical ($h$), and horizontal ($w$) axes in a fine-grained, round-robin manner, as shown in Figure 3. This design ensures that each positional axis is encoded using the full frequency spectrum, from high to low, thereby enabling robust multi-scale modeling for each positional axis. Moreover, the uniform frequency distribution of our interleaved design is compatible with extrapolation algorithms like NTK-aware (bloc97, 2023) and YaRN (Peng et al., 2024), which function by rescaling the frequency spectrum (see Appendix D.3).
@@
From the Position Design aspect, methods like IL-RoPE (Liao et al., 2025) and Omni-RoPE (Wu et al., 2025) modify the text encoding. As shown in Figure 1(e), they reset spatial coordinates for each image to aid editing but concurrently set the spatial dimensions for text tokens to zero. This design choice breaks compatibility with the standard RoPE used in the pre-trained LLM.
@@ 3.1 Experimental Setup
Training Data and Evaluation Benchmarks. We conduct experiments using approximately 2M high-quality supervised fine-tuning (SFT) samples, covering a wide range of visual scenarios including image captioning, OCR, visual reasoning, visual grounding, document comprehension, and long video understanding. For evaluation, we adopt more than 20 benchmarks spanning images, videos, and grounding tasks. Specifically, the image benchmarks include MMMU (Yue et al., 2024), MMBench (Liu et al., 2024a), MMStar (Chen et al., 2024), OCRBench (Liu et al., 2024b), AI2D (Kembhavi et al., 2016), RealWorldQA (X.AI., 2024), DocVQA (Mathew et al., 2021), TextVQA Singh et al. (2019), InfoVQA (Mathew et al., 2022), and ChartQA (Masry et al., 2022). The video benchmarks consist of MVBench (Li et al., 2024), STAR (Wu et al., 2021), VideoMME (Fu et al., 2025), LVBench (Wang et al., 2024b), MLVU (Zhou et al., 2024b), and Charades-STA (Zhou et al., 2024a). For grounding, we evaluate on RefCOCO (Kazemzadeh et al., 2014) series.
@@
Types Benchmarks Vanilla RoPE MRoPE VideoRoPE HoPE CircleRoPE MHRoPE MRoPE-I Image MMMU 50.56 50.22 49.89 49.89 47.22 53.00 53.22(+2.67) $\text{MMBench}_{avg}$ 74.75 74.06 75.95 75.35 74.91 75.04 75.56(+0.81) MMstar 49.13 49.93 49.60 50.33 47.00 49.60 51.13(+2.00) OCRBench 73.40 72.70 66.20 66.60 70.60 73.40 74.00(+0.50) AI2D 76.20 74.94 74.29 76.10 74.45 75.45 75.36 (-0.84) RealworldQA 58.30 57.25 56.21 57.12 56.60 60.52 57.39 (-0.92) DocVQA 82.94 81.49 60.13 60.12 77.70 81.32 83.72(+0.78) TextVQA 66.80 65.85 66.58 66.77 65.54 66.49 66.91(+0.11) InfoVQA 58.85 52.96 37.42 34.80 53.11 52.01 58.24(-0.62) ChartQA 56.84 63.56 54.88 55.44 53.72 62.44 62.12 (+5.28) Video MVBench 57.05 57.85 56.78 58.00 57.10 58.93 57.05 (+0.00) STAR 58.07 58.28 57.20 58.30 57.94 59.48 57.79 (-0.28) MLVU 64.69 63.26 66.05 64.81 62.69 65.69 65.46 (+0.77) VideoMME 58.63 58.22 58.70 59.52 57.70 57.48 58.96(+0.33) LVBench 38.93 39.22 40.15 40.99 38.80 40.32 40.54(+1.61) Charades-STA 32.49 32.23 34.21 36.07 32.27 33.56 34.36(+1.87) Grounding $\text{RefCOCO}_{val}$ 77.67 78.35 77.95 77.72 79.59 79.87 80.94(+3.27) un $\text{RefCOCO}_{testA}$ 81.37 82.52 80.43 81.60 83.98 83.66 84.55(+3.18) $\text{RefCOCO}_{testB}$ 72.66 72.31 72.62 71.44 74.35 73.20 75.05(+2.39) $\text{RefCOCO+}_{val}$ 69.16 68.80 68.15 69.61 70.19 70.55 71.80(+2.64) $\text{RefCOCO+}_{testA}$ 74.48 75.95 73.14 74.55 76.77 76.88 77.44(+2.96) $\text{RefCOCO+}_{testB}$ 61.67 59.97 60.50 61.69 62.59 61.96 61.96(+0.29) $\text{RefCOCOg}_{val}$ 75.45 75.86 74.06 74.69 76.10 76.55 77.70(+2.25) $\text{RefCOCOg}_{test}$ 75.40 75.73 73.90 75.45 76.12 76.68 77.34(+1.94) Overall Image 65.69 65.18 60.64 60.72 62.86 66.40 66.65(+0.96) Video 51.64 51.51 52.18 52.95 51.09 52.98 52.36(+0.72) Grounding 73.48 73.69 72.59 73.19 74.96 74.92 75.85(+2.36)
@@ 3.3.1 Ablation Study on Position Design
Enlarged Modality Interval: We also tested artificially enlarging the modality interval to match that of vanilla RoPE, a strategy similar to RoPE-Tie (Su, 2024). This also resulted in poor document-related performance. However, the failure mode was distinct: the model generated fluent but contextually irrelevant text, effectively ignoring the visual input. This suggests that while a clear modality interval is necessary, simply maximizing its size to align with vanilla RoPE can be detrimental by
@@ D.3 Compatibility with YaRN Extrapolation
As shown in Figure 4(b), the interleaved frequency allocation of MRoPE-I makes it compatible with extrapolation algorithms like NTK-aware (bloc97, 2023) and YaRN (Peng et al., 2024). Whereas standard MRoPE’s partitioned spectrum complicates the application of a consistent frequency scaling boundary, our interleaved design provides a full spectrum across all positional axes, enabling a straightforward and symmetric application of these methods.
@@ D.4 Long-Context Video Understanding
We further compare the performance of different methods on long video understanding, with context lengths ranging from 32K to 256K. As shown in Figure 5, apart from LVBench, we do not observe clear performance improvements or degradation when extrapolating to longer sequences. The only exception is Vanilla RoPE, which suffers from a sharp performance drop at 128K/256K. We attribute this to excessively fast-growing position IDs, which lead to degraded extrapolation capability, which also discussed in other works (Wei et al., 2025; Li et al., 2025).

### v2

@@ 1 Introduction
The permutation-invariant nature of the self-attention mechanism requires the use of positional encodings to inform Large Language Models (LLMs) of sequence order, relative distance, and structural dependencies. While early methods relied on absolute position embeddings (transformer), relative encodings—which better generalize to varying sequence lengths—have become the standard. Among these, Rotary Position Embedding (RoPE) (roformer) has emerged as the de facto choice in modern LLMs such as Llama (llama3) and Qwen (qwen3).
Vision-Language Models (VLMs) also require positional encodings that can handle heterogeneous modalities, including 1D text and 2D/3D visual inputs. Current methods fall into two main categories: 1D sequential and multi-dimensional designs. The former, exemplified by vanilla RoPE (roformer) and V2PE (v2pe), flattens and concatenates all inputs into a single sequence. While simple, this approach discards the native visual geometry, leading to a significant degradation in performance on tasks requiring visual grounding and spatial reasoning.
Multi-dimensional designs, the second approach, extend RoPE to multiple axes (time, height, width) by partitioning embedding channels. Qwen2‑VL qwen2vl adopts Multimodal RoPE (MRoPE) to unify positional encoding for text and visual tokens. However, MRoPE allocates the position embedding into t-h-w chunk, placing the temporal infomation entirely in the high‑frequency channels. This bias in temporal encoding harms long‑range video modeling. Subsequent work has attempted to improve it, but this has led to a fragmented landscape of highly specialized solutions. Some methods focus exclusively on image understanding (circlerope), others on video comprehension (videorope; hope; vrope), and a third group on image generation (ilrope; omnirope). While these models achieve notable performance in their respective domains, the development of a truly robust and versatile VLM requires a more holistic positional encoding strategy. In this work, we aim to develop a more holistic positional encoding strategy capable of supporting the core, unified capabilities of image and video understanding, complemented by fine-grained visual grounding.
@@ 2.1 Preliminaries: Vanilla RoPE
Vanilla RoPE (roformer) is a pivotal method for encoding positional information in modern LLMs. Unlike additive position embeddings, RoPE applies a rotational transformation to the query and key vectors, thereby incorporating relative position dependencies directly into the self-attention mechanism. Given a query vector $\bm{q}$ at position $m$ and a key vector $\bm{k}$ at position $n$, the attention scores $\bm{S}$ are calculated as:
@@ 2.2 Position Design
The most straightforward approach, employed by vanilla RoPE (roformer) and V2PE (v2pe), is to treat the multimodal input as a flattened, one-dimensional sequence. Position indices are assigned incrementally, with the position $m_{i}$ of the $i$-th token defined as $m_{i}=m_{i-1}+s_{\text{mod}}$, where $s_{\text{mod}}$ is a step size specific to the token’s modality. For vanilla RoPE, all modalities are treated uniformly, with $s=1$.
As shown in Figure 1(a), this design presents two significant drawbacks. First, it discards the inherent 3D structure of visual content, which can alter the spatio-temporal reasoning capabilities of a VLM. Second, position indices can grow exceedingly large in long sequences, negatively affecting the model’s extrapolation performance (videorope).
To address the issue of large position indices, V2PE (v2pe) introduces dynamic position scaling for visual tokens, setting their step size $s_{\text{visual}}$ to a value in $\{1,1/2,\dots,1/256\}$. This modification mitigates the rapid growth of position indices and has shown benefits in long video understanding. However, the 3D structure of visual content is still ignored in 1D sequential design.
@@
To preserve the native 3D structure of visual content, methods like MRoPE (qwen2vl) (See Figure 1(b)) extend the scalar position identifier to a multi-dimensional tuple. For instance, a token’s position can be represented as $\bm{m}_{i}=(m_{i}^{t},m_{i}^{h},m_{i}^{w})$, corresponding to its temporal, vertical, and horizontal axes. MRoPE conceptually treats each visual content (e.g., an image or a set of video frames) as a single, large ”cube”. The temporal position of the subsequent token is then set by ”jumping” past the maximum coordinate value of current block. This is achieved with the update rule:
@@
This strategy guarantees that no positional overlaps occur between modalities. However, proponents of VideoRoPE (hope) and HoPE (hope) argue that MRoPE’s position design lacks inter-modal symmetry, they introduce a ”diagonal layout” by centering the spatial coordinates (see Figure 1(c)). In this scheme, visual frames are not only stacked along the temporal axis but are also shifted along the vertical and horizontal axes. Despite its theoretical elegance, this diagonal layout introduces a critical flaw: the potential for position id overlap between visual content and generated text tokens. For high-resolution image content like documents, the spatial coordinates of visual tokens can extend into the index range subsequently assigned to the generated text tokens. We identify this positional ambiguity as a source of ”modalities confusion in generation”, a failure mode that manifested as endless text repetition in our later experiments.
CircleRoPE (circlerope) arranges image tokens in a circular layout, orthogonal to the linear axis of text positions (see Figure 1(d)). A key property of this design is that it renders all visual tokens equidistant from any given text token, which theoretically promotes uniform attention across the image. However, CircleRoPE’s design has two limitations. First, the large interval between modalities may impede effective cross-modal interaction. Second, lacking a temporal axis, it collapses all video frames onto a single ring, which introduces severe temporal ambiguity.
@@ 2.3.3 Towards an Optimal Frequency Allocation
Multi-Head Allocation. The first strategy, termed Multi-Head RoPE (MHRoPE), is inspired by recent work demonstrating channel-level redundancy in RoPE (e.g., partial RoPE (partialrope)). Based on the premise that similar redundancy exists at the attention head level, MHRoPE partitions the positional encoding task among different attention heads [^4] For Group Query Attention (GQA), we partition on KV heads and repeat on corresponding query heads., as shown in Figure 3. A primary advantage of this strategy is that each axis is encoded using the full frequency spectrum available within its assigned heads. This approach avoids the loss of frequency resolution inherent to channel-splitting methods. Moreover, it may be more scalable. As the number of positional axes grows (vrope), partitioning a fixed channel channels (e.g., 128) becomes untenable, whereas dedicating distinct heads to new dimensions offers a far more robust and flexible approach.
Interleaved Allocation. Our second strategy, employed in our MRoPE-Interleaved (MRoPE-I) method, distributes feature channels to the temporal ($t$), vertical ($h$), and horizontal ($w$) axes in a fine-grained, round-robin manner, as shown in Figure 3. This design ensures that each positional axis is encoded using the full frequency spectrum, from high to low, thereby enabling robust multi-scale modeling for each positional axis. Moreover, the uniform frequency distribution of our interleaved design is compatible with extrapolation algorithms like NTK-aware (ntk-aware) and YaRN (yarn), which function by rescaling the frequency spectrum (see Appendix D.3).
@@
From the Position Design aspect, methods like IL-RoPE (ilrope) and Omni-RoPE (omnirope) modify the text encoding. As shown in Figure 1(e), they reset spatial coordinates for each image to aid editing but concurrently set the spatial dimensions for text tokens to zero. This design choice breaks compatibility with the standard RoPE used in the pre-trained LLM.
@@ 3.1 Experimental Setup
Training Data and Evaluation Benchmarks. We conduct experiments using approximately 2M high-quality supervised fine-tuning (SFT) samples, covering a wide range of visual scenarios including image captioning, OCR, visual reasoning, visual grounding, document comprehension, and long video understanding. For evaluation, we adopt more than 20 benchmarks spanning images, videos, and grounding tasks. Specifically, the image benchmarks include MMMU (mmmu), MMBench (mmbench), MMStar (mmstar), OCRBench (ocrbench), AI2D (ai2d), RealWorldQA (realworldqa), DocVQA (docvqa), TextVQA textvqa, InfoVQA (infovqa), and ChartQA (chartqa). The video benchmarks consist of MVBench (mvbench), STAR (star), VideoMME (videomme), LVBench (lvbench), MLVU (mlvu), and Charades-STA (charades-sta). For grounding, we evaluate on RefCOCO (refcoco) series.
@@
Types Benchmarks Vanilla RoPE MRoPE VideoRoPE HoPE CircleRoPE MHRoPE MRoPE-I Image MMMU 50.56 50.22 49.89 49.89 47.22 53.00 53.22(+2.67) $\text{MMBench}_{avg}$ 74.75 74.06 75.95 75.35 74.91 75.04 75.56(+0.81) MMstar 49.13 49.93 49.60 50.33 47.00 49.60 51.13(+2.00) OCRBench 73.40 72.70 66.20 66.60 70.60 73.40 74.00(+0.50) AI2D 76.20 74.94 74.29 76.10 74.45 75.45 75.36 (-0.84) RealworldQA 58.30 57.25 56.21 57.12 56.60 60.52 57.39 (-0.92) DocVQA 82.94 81.49 60.13 60.12 77.70 81.32 83.72(+0.78) TextVQA 66.80 65.85 66.58 66.77 65.54 66.49 66.91(+0.11) InfoVQA 58.85 52.96 37.42 34.80 53.11 52.01 58.24(-0.62) ChartQA 56.84 63.56 54.88 55.44 53.72 62.44 62.12 (+5.28) Video MVBench 57.05 57.85 56.78 58.00 57.10 58.93 57.05 (+0.00) STAR 58.07 58.28 57.20 58.30 57.94 59.48 57.79 (-0.28) MLVU 64.69 63.26 66.05 64.81 62.69 65.69 65.46 (+0.77) VideoMME 58.63 58.22 58.70 59.52 57.70 57.48 58.96(+0.33) LVBench 38.93 39.22 40.15 40.99 38.80 40.32 40.54(+1.61) Charades-STA 32.49 32.23 34.21 36.07 32.27 33.56 34.36(+1.87) Grounding $\text{RefCOCO}_{val}$ 77.67 78.35 77.95 77.72 79.59 79.87 80.94(+3.27) $\text{RefCOCO}_{testA}$ 81.37 82.52 80.43 81.60 83.98 83.66 84.55(+3.18) $\text{RefCOCO}_{testB}$ 72.66 72.31 72.62 71.44 74.35 73.20 75.05(+2.39) $\text{RefCOCO+}_{val}$ 69.16 68.80 68.15 69.61 70.19 70.55 71.80(+2.64) $\text{RefCOCO+}_{testA}$ 74.48 75.95 73.14 74.55 76.77 76.88 77.44(+2.96) $\text{RefCOCO+}_{testB}$ 61.67 59.97 60.50 61.69 62.59 61.96 61.96(+0.29) $\text{RefCOCOg}_{val}$ 75.45 75.86 74.06 74.69 76.10 76.55 77.70(+2.25) $\text{RefCOCOg}_{test}$ 75.40 75.73 73.90 75.45 76.12 76.68 77.34(+1.94) Overall Image 65.69 65.18 60.64 60.72 62.86 66.40 66.65(+0.96) Video 51.64 51.51 52.18 52.95 51.09 52.98 52.36(+0.72) Grounding 73.48 73.69 72.59 73.19 74.96 74.92 75.85(+2.36)
@@ 3.3.1 Ablation Study on Position Design
Enlarged Modality Interval: We also tested artificially enlarging the modality interval to match that of vanilla RoPE, a strategy similar to RoPE-Tie (ropetie). This also resulted in poor document-related performance. However, the failure mode was distinct: the model generated fluent but contextually irrelevant text, effectively ignoring the visual input. This suggests that while a clear modality interval is necessary, simply maximizing its size to align with vanilla RoPE can be detrimental by
@@ D.3 Compatibility with YaRN Extrapolation
As shown in Figure 4(b), the interleaved frequency allocation of MRoPE-I makes it compatible with extrapolation algorithms like NTK-aware (ntk-aware) and YaRN (yarn). Whereas standard MRoPE’s partitioned spectrum complicates the application of a consistent frequency scaling boundary, our interleaved design provides a full spectrum across all positional axes, enabling a straightforward and symmetric application of these methods.
@@ D.4 Long-Context Video Understanding
We further compare the performance of different methods on long video understanding, with context lengths ranging from 32K to 256K. As shown in Figure 5, apart from LVBench, we do not observe clear performance improvements or degradation when extrapolating to longer sequences. The only exception is Vanilla RoPE, which suffers from a sharp performance drop at 128K/256K. We attribute this to excessively fast-growing position IDs, which lead to degraded extrapolation capability, which also discussed in other works (videorope; hope).